a rectangular parking lot has the length of 7 yards graeter than the width. the area is 330 square yards. find the length and width?
Answer by mananth(12270)

let width be x
length = x+7
x(x+7)=330
x^2+7x-330=0
x^2+22x-15x-330 =0
x(x=22)-15(x=22)=0
(x-22)(x-15)=0
x=-22 , x= 15 ignore negative value of 22
The width is 15 yards.
the length will be 15+7 = 22 yards
..
CHECK
15*22=330 sq. yards